# Anamorphic catadioptric telescope and generalized spherical aberration

### v3

@@ 3.2 Doublet lens set (II)
Since two refractive surfaces with vanishing conic constants is not enough, we consider the set (II) containing three optical surfaces. Specifically, we choose an anamorphic extension of a refractive doublet objective, contact type, in its standard form with positive front lens and negative rare one, i.e. we take two connected toroidal lenses, positive and negative, with circular YZ profiles. We denote the refractive indices of the positive and negative components by $n$ and $\tilde{n}$, respectively, and chose $n<\tilde{n}$. The respective condition for compensation of GS aberrations reads
@@
The Eq. (29) nearly vanishes for a broad range of values of the parameters.
We remark that the used method can be applied to more complex refractive sets and extended to compensate for other aberrations besides the spherical ones.
@@ 3.3 Proof of concept by image simulation
To validate the proposed concept design we consider f/10 anamorphic telescope with two-mirror set (I) and doublet lens set (II), which has two equal focal ratios:
@@
where $L_{x}$ and $L_{y}$ are the sides of its rectangular aperture. To be more specific, we choose the set (I), which is a cylindrical analog of RC-type Cassegrain. For a set (II) we take toroidal analog of the primary spherical aberration and coma free apochromatic doublet [22], since the discussed above economical three-surface contact doublet generates several times larger RMS spot.
@@
We simulated this OS and performed ray-tracing using the optical design software CODE V [23]. The simulation is performed for a on-axis $0^{\circ}$ field point and the wavelength 587.6 nm. The set (I) parameter values we take as follows: $F_{1}=1000$ mm, $R_{1}=450$ mm, $R_{2}=210$ mm, $\kappa_{1}=-1.057$ and $\kappa_{2}=-2.839$. Adopting the apochromatic doublet specification [22] to provide the ratio of foci of $F_{1}/F_{2}\approx 5$, for the set (II) we obtain the following set of curvature radii with the sign factors {$95.75$ mm, $-41.98$ mm, $-41.98$ mm, $-367.74$ mm} and the lens separation of 0.688 mm, where the positive and negative components widths are $d_{L}=3.704$ mm and $\tilde{d}_{L}=2$ mm, and their glass types are FPL52 (OHARA) and ZKN7 (SCHOTT), respectively. The XZ, YZ and prospective views of the anamorphic telescope are shown in Fig. 5.
Figure 5: From top to bottom: XZ view, YZ view and prospective view for the considered realization of an anamorphic catadioptric telescope with the RC-type Cassegrain set (I) and the apochromatic doublet set (II).
Figure 6: Spot diagrams for the discussed anamorphic telescope at the image plane (left) and best focus (right).
The spot diagrams for a on-axis $0^{\circ}$ field point are shown in Fig. 6, where the circle represents a contour of Airy disk, which size is $2.44\times\lambda\times 10$ with 10 standing for the focal ratio. The RMS spot size is 1.5$\pm$0.1 $\mu$m at the image plane and 1.25$\pm$0.1 $\mu$m at the best focus. Clearly, the shown spot diagrams possess double-plane symmetry just like the OS itself.
Next we compare performances of the axially symmetric analogs of isolated sets (I) and (II). The prospective view and spot diagram for such analog of set (I) with the quadratic aperture $L_{x}\times L_{x}$ is shown in Fig. 7. Its RMS spot size is only 0.06$\pm$0.1 $\mu$m. The spot diagrams for the analog of set (II) with the quadratic aperture $L_{y}\times L_{y}$ is shown in Fig. 8. Its RMS spot size is 3.3$\pm$0.1 $\mu$m at the image plane and 1.19$\pm$0.1 $\mu$m at the best focus. For plotting spot diagrams we used the grid ray pattern with 20 rays across the aperture. It is clear from the spot diagrams in Figs. 7 (right) and 8 that aberrations of the considered anamorphic catadioptric telescope are dominated by the aberrations of its refractive set (II).
@@
Figure 7: The prospective view (left) and spot diagram (right) for the axially symmetric analog of isolated set (I) with the aperture 100 mm $\times$ 100 mm.
@@
Figure 8: Spot diagrams at the image plane (left) and best focus (right) for the axially symmetric analog of isolated set (II) with the aperture 20 mm $\times$ 20 mm.
@@ 3.4 Field of view of the anamorphic telescope
Let’s compare field of view (FOV) of an anamorphic telescope (AT) possessing the two focal distances $F_{1}$ and $F_{2}$, where $F_{1}>F_{2}$, and FOV of usual rotationally symmetric telescope (RST) with the focal distance $F_{1}$. Suppose that they are equipped with the identical eyepieces, which are specified by the focal distance $f$ and quadratic field stop, i.e. FOV $\varphi\times\varphi$. Then FOV of RST is equal angle:
@@
We conclude that AT possesses $F_{1}/F_{2}$ times wider FOV and uses its CCD camera $F_{1}/F_{2}$ times more effectively with respect to RST of the same rectangular aperture (assuming that CCD camera is a standard quadratic photosensor array).
@@ 4 Discussion and conclusion
Following a brief intro on anamorphic OS aberrations and their possible partial compensations, we demonstrated its application to design of an anamorphic telescope, in particular, for GS aberration compensation in the OS symmetry planes. We proposed an anamorphic catadioptric telescope that combines a translationally invariant reflective part and an anamorphic refractive part, which has a rotational symmetry with respect to the axis that is parallel to the direction of translational symmetry of the reflective part and passes through the OS focal point. We outlined derivation of analytical expressions for the GS aberration compensations discussed, considering OS with cylindrical version of the Cassegrain reflector as the reflective part and two examples of the refractive part, which are relatively simple for analytical consideration. Then we simulated the anamorphic telescope combining the cylindrical RC-type Cassegrain with toroidal apochromatic doublet, using the ray-tracing software CODE V [23]. This simulation demonstrated good OS performance, which can be further improved by using more advanced refractive set and introducing corrective lenses.
In general, a wide class of cylindrical versions of known telescopic objectives, either reflective or refractive, can serve as a set (I) and variety of toroidal versions of known axially symmetric refractive telescopic OS can serve as a set (II). (We remark that reflective anamorphic set (II) is also possible [DZ, work in progress].) One of the main limitations comes from avoiding light blocking by the set (II), e.g., the distance to the focus behind the secondary mirror must be sufficiently greater than the focal distance $F_{2}$ of the set (II). The expressions for aberration compensation for the OS symmetry planes XZ and YZ (in the discussed coordinates) are identical to the ones for axially symmetric analogs of the sets (I) and (II), respectively. New analytical expressions can be found for aberration compensation beyond these planes, which is an important direction of future research.
@@
To conclude, we presented a viable concept of an anamorphic telescope of a new type that can possess two large magnifications in the orthogonal directions at the image plane corresponding to the two independent focal distances, which allow to control both dimensions of the filed of view.
@@ Acknowledgments
The author thanks Barbara Pintaske and Daniela Ponce for their attention when providing the optical software of the SYNOPSYS.

### v4

@@ 3.2 Doublet lens set (II)
Since two refractive surfaces with vanishing conic constants is not enough to make the set (II) free of GS aberration, we consider its version containing three optical surfaces. Specifically, we choose an anamorphic extension of a refractive doublet objective, contact type, in its standard form with positive front lens and negative rare one, i.e. we take two connected toroidal lenses, positive and negative, with circular YZ profiles. We denote the refractive indices of the positive and negative components by $n$ and $\tilde{n}$, respectively, and chose $n<\tilde{n}$. The respective condition for compensation of GS aberrations reads
@@
The Eq. (30) nearly vanishes for a broad range of values for the six parameters, which are essential in this consideration: $r_{3}$, $r_{4}$, $d_{L}$, $\tilde{d}_{L}$, $n$ and $\tilde{n}$.
We remark that the used method can be applied to more complex refractive sets and extended to compensate for other aberrations besides the spherical one.
Figure 7: XZ view (top) and YZ view (bottom) of the considered realization of an anamorphic catadioptric telescope with the RC-type Cassegrain set (I) and the apochromatic doublet set (II).
Figure 8: Perspective view of the considered realization of an anamorphic catadioptric telescope with the RC-type Cassegrain set (I) and the apochromatic doublet set (II).
Figure 9: Spot diagram for the discussed anamorphic telescope at the image plane for the specified five fields and three focal positions.
Figure 10: OPD plot for the discussed anamorphic telescope for the specified five fields.
@@ 3.3 Proof of concept by image simulation
To validate the proposed concept design we consider f/10 anamorphic telescope (AT) with two-mirror set (I) and doublet lens set (II), which has two equal focal ratios:
@@
where $L_{x}$ and $L_{y}$ are the sides of its rectangular aperture. To be more specific, we choose the set (I), which is a cylindrical analog of a Ritchey-Chretien type Cassegrain. For a set (II) we take toroidal analog of the primary spherical aberration and coma free apochromatic doublet [22], since the simplified three-surface contact doublet, which is discussed in section 3.2, generates essentially larger aberrations. Having the reflective set (I) and apochromatic set (II), we obtain OS that is nearly free of chromatic aberrations. We notice that design optimization for specific narrow observation targets requires a deeper research on selection of a particular set (II).
@@
We simulated this OS and performed ray-tracing using the optical design software CODE V [23]. The simulation is performed for three wavelengths: d, F, and C (the spectral lines for helium at 587.6 nm and for hydrogen at 486.1 and 656.3 nm, respectively). The set (I) parameter values we take as follows: $F_{1}=1000$ mm, $R_{1}=450$ mm, $R_{2}=210$ mm, $\kappa_{1}=-1.057$ and $\kappa_{2}=-2.839$. Adopting the apochromatic doublet specification [22] to provide the ratio of foci of $F_{1}/F_{2}\approx 5$, and optimizing consequently the radii of curvature $r_{5}$ and $r_{3}$, for the set (II) we obtain the following set of curvature radii from $r_{3}$ to $r_{6}$ with the sign factors {$95.75$ mm, $-41.98$ mm, $-41.98$ mm, $-367.74$ mm} and the lens separation of 0.688 mm, where the positive and negative components’ widths are $d_{L}=3.704$ mm and $\tilde{d}_{L}=2$ mm, and their glass types are FPL52 (OHARA) and ZKN7 (SCHOTT), respectively. In the above optimization we applied the standard CODE V procedure with the ‘Transverse Ray Aberration’ error function. In result, performance of the set (II) was improved in both modes of isolation and as part of the AT.
The XZ, YZ and perspective views of the resulting AT are shown in Figs. 7 and 8. Its spot diagram is drawn in Fig. 9 for three defined focal positions and five chosen field points: on-axis $0^{\circ}$ and two off-axis pairs in the XZ and YZ planes. The Y field angles are taken $F_{1}/F_{2}$ times larger then the X field angles (this is natural according to section 3.4). This spot diagram and the following ones are generated by tracing an evenly spaced rectangular grid of rays on the entrance pupil from each wavelength with 51 rays across the aperture used for plotting. The spot diagrams list, for each field point and zoom position, the following: the relative field heights with X and Y field angles in the left, and the minimum RMS (with the fractional error of $10^{-4}$) and 100% spot size diameters in the right. The corresponding optical path difference (OPD) plot is shown in Fig. 10. From Figs. 9 and 10 it is clear that the proposed AT can serve only as a narrow field of view instrument.
@@
Figure 11: The perspective view of the axially symmetric analog of isolated set (I) with the aperture 100 mm $\times$ 100 mm.
@@
Figure 12: The spot diagram for the axially symmetric analog of isolated set (I) with the aperture 100 mm $\times$ 100 mm.
Figure 13: The OPD plot for the axially symmetric analog of isolated set (I) with the aperture 100 mm $\times$ 100 mm.
Figure 14: The perspective view of the axially symmetric analog of isolated set (II) with the aperture 20 mm $\times$ 20 mm.
Figure 15: Spot diagram at the image plane for the axially symmetric analog of isolated set (II) with the aperture 20 mm $\times$ 20 mm.
Figure 16: The OPD plot for the axially symmetric analog of isolated set (II) with the aperture 20 mm $\times$ 20 mm.
Next we analyze performances of the axially symmetric analogs of isolated sets (I) and (II), which perform similarly to the original sets in the XZ and YZ planes, correspondingly. At the same time these modified sets possess focal points that allows for straightforward comparison with the AT performance by means of spot diagram and OPD plot. The perspective view, spot diagram and OPD plot for such analog of set (I) with quadratic $L_{x}\times L_{x}$ aperture are shown in Figs. 11, 12 and 13, respectively. The perspective view, spot diagram and the OPD plot for the analog of set (II) with quadratic $L_{y}\times L_{y}$ aperture are shown in Figs. 14, 15 and 16, correspondingly. The black circumferences in Fig. 14 represent contours of Airy disk, which size is $2.44\times\lambda\times 10$ with 10 standing for the focal ratio. Next we track the effect of rectangular pupil shape on performance of the set (I). The perspective view, spot diagram and OPD plot for the axially symmetric analog of set (I) with rectangular $L_{x}\times L_{y}$ aperture are shown, respectively, in Figs. 17, 18 and 19. It is clear from the above spot diagrams and OPD plots that for near-axis field points aberrations of the considered AT are mainly originated from its refractive set (II). However, for the object angle of 0.3 and higher the aberrations associated with the reflective set (I) become more essential. We pay readers’ attention on the different scale for the upper entries in Figs. 12 and 18, and notice drastic increase of a spot size with the object angle increase for an isolated the set (I). The introduction of the set (II) allows us to essentially suppress this effect and reduce X component of transverse aberration for the AT, and decisively improve its image quality.
We remark that spot diagram of the anamorphic analog of isolated set (I) in Fig. 18 possesses two mirror symmetries with respect to the axes X and Y. However, spot diagram of the AT in Fig. 9 conserves only one mirror symmetry even for slightly off-axial fields in the XZ plane.
Figure 17: The perspective view of the axially symmetric analog of isolated set (I) with the rectangular aperture 100 mm $\times$ 20 mm.
Figure 18: The spot diagram for the axially symmetric analog of isolated set (I) with the rectangular aperture 100 mm $\times$ 20 mm.
Figure 19: The OPD plot for the axially symmetric analog of isolated set (I) with the rectangular aperture 100 mm $\times$ 20 mm.
Figure 20: The spot diagrams for the anamorphic telescope (upper) and the axially symmetric analog of its isolated set (I) (lower), for the same rectangular aperture and plotting scale.
@@ 3.4 Field of view of the anamorphic telescope
Let’s compare field of view (FOV) of a usual rotationally symmetric telescope (RST) with the focal distance $F_{1}$ and FOV of the discussed AT possessing the two focal distances $F_{1}$ and $F_{2}$, where $F_{1}$ is larger than $F_{2}$. Suppose that they are equipped with the identical eyepieces, which are specified by the focal distance $f$ and quadratic field stop, i.e. apparent FOV $\varphi\times\varphi$. Then FOV of RST is equal angle:
@@
We conclude that AT possesses $F_{1}/F_{2}$ times wider FOV in the Y direction with respect to RST of the same rectangular aperture. In particular, in the above examples of the AT (Figs. 7 and 8) and the RST (Fig. 17), which are identical in size, the former has five times wider FOV in the Y direction with respect to the later if both are equipped with identical eyepieces. We remark that in case of RST essential increase of FOV in the Y direction results in dramatic growth of aberrations that is shown in Fig. 20, which compares the spot diagrams for AT and RST for the same two fields with nonzero Y object angles. Similarly, in case of using a rectangular CCD camera or active-pixel sensor, AT provides $F_{1}/F_{2}$ times wider FOV in the Y direction, along which magnification is $F_{1}/F_{2}$ times lower.
@@ 4 Discussion and conclusion
After a brief intro on anamorphic OS aberrations and their partial compensation, we presented the following application. We proposed an anamorphic catadioptric telescope that combines translationally invariant reflective part and an anamorphic refractive part, which has a rotational symmetry about the focal segment of the reflective part. We outlined derivation of analytical expressions for the GS aberration compensations in the OS symmetry planes, in the case of cylindrical version of the Cassegrain reflector as the reflective part and two alternatives for the refractive part, which are relatively simple for analytical consideration. Then we simulated the anamorphic telescope combining the cylindrical RC-type Cassegrain with toroidal apochromatic doublet, using the ray-tracing software CODE V [23]. This simulation demonstrated good OS performance, which can be further improved by using more advanced refractive set and introducing corrective lenses.
In general, a wide class of cylindrical versions of known telescopic objectives, reflective, refractive or catadioptric, can serve as a set (I) and variety of generalized toroidal versions of known axially symmetric reflective, refractive or catadioptric telescopic OS can serve as a set (II) (DZ, patent submission Nr. P.452264 to the Patent Office of the Republic of Poland, 2025). Moreover, the presented method of separate focusing in the two dimensions can be extended to off-axis systems.
One of possible limitations for the introduced anamorphic OS comes from avoiding light blocking by the components of set (II), e.g., the distance to the focus behind the secondary mirror must be sufficiently greater than the focal distance $F_{2}$ of the set (II). However, this issue is not crucial for a wide class of solutions, in particular, such blocking does not take place for the examples shown in section 3.3. The expressions for aberration compensation for the OS symmetry planes XZ and YZ (in the discussed coordinates) are identical to the ones for axially symmetric analogs of the sets (I) and (II), respectively. Aberration compensation beyond these planes, including chromatic and field curvature aberrations, is an important direction of future research. Several simplifications leave us with a simple two-step design procedure for the refractive set (II): (a) all parameters for the YZ plane (refractive indices, radii of curvature, lens thicknesses and separations, etc.) can be derived using the same methods as for ordinary rotationally symmetric OS, (b) each optical element of the set (II) can be obtained as the body of rotation of a given YZ profile with respect to the focal line of the set (I) by a proper angle to collect the converging light beam from the set (I).
The introduced optical scheme can be utilized in the design of terrestrial and space telescopes of various functional purposes, which would require inclusion of FOV and optimization to correct for coma, etc. However, the goal of this paper is a concept presentation of a new OS class rather than a complete design of its particular realization.
Obvious benefits of the discussed anamorphic telescope, such as weight saving and fittability, arise from the shape of its primary mirror. A rectangular shape of a primary in the XY projection can provide economy in mass for the same resolution in the X (larger side) direction compared to a circular shape, while rotation of the telescope along the Z axis governs direction of its best resolution. The size of mass saving can be easily estimated. Assuming that mass density and thickness of the primaries of a rotationally symmetric telescope and an anamorphic one are the same and constant, the ratio of masses of these mirrors, which essentially contribute to the masses of the instruments, is equal to the ratio of their surface areas, $M_{\rm RST}/M_{\rm AT}=S_{\rm RST}/S_{\rm AT}$. Taking the diameter of the RST’s circular aperture to be equal to the larger side $L_{x}$ of the AT’s rectangular aperture $L_{x}\times L_{y}$, for the case of circular $XZ$ profiles the ratio of their surface areas can be written as
$\frac{S_{\rm RST}}{S_{\rm AT}}=\frac{\pi h}{L_{y}\arccos\left(\frac{R-h}{R}\right)},$ (39)
where $h$ is the depth (sag) and $R$ is the radius of curvature in the XZ plane. (Strictly speaking, one has to compare instruments of equal maximal resolution rather then equal maximal aperture size [15]. Modification of the Rayleigh criterion for the anamorphic OS, which we ignore here, may further increase the discussed gain in mass.) Using the relation
$L_{x}=2\sqrt{2Rh-h^{2}},$ (40)
for small $h$ with respect to $R$ we obtain an estimate for the ratio of masses of the primaries in the the following simple form
$\frac{M_{\rm RST}}{M_{\rm AT}}\approx\frac{\pi}{4}\frac{L_{x}}{L_{y}},$ (41)
which is equal to the ratio of respective aperture areas as expected. A more narrow AT’s primary provides larger gain in mass. However, a more distorted image is generated that complicates reconstruction of the object’s appearance.
Potentially, the discussed two-step focusing scheme can be useful also when designing solar observatories. Additionally, reduction of manufacturing costs is possible for a cylindrical primary (in particular, produced by bending a flat mirror surface) compared to a rotationally symmetric one. However, adaptation of polishing technology and testing methods is required.
One of more traditional approaches to weight saving without loss of aperture size is in taking a narrow slice of a rotationally symmetric primary. This method can significantly reduce costs in case of large but stationary aperture. An interesting solution is RATAN-600 telescope, which primary comprises a 576 m diameter circle of rectangular radio reflectors that can be angled causing the overall effect of a partially steerable antenna. This design works well for the radio waves. However, it is hard to apply it for much shorter wavelengths that would require much higher accuracy in positioning and shaping the reflecting parts. Concerning possible space applications, this approach can benefit in weight saving, but would suffer from dimensional restrictions existing for payloads. On the contrary, the introduced anamorphic solution allows for fully steerable ground-based telescopes with large aperture in one dimension. Moreover, in space applications, it reconciles large aperture (maximal) size, large enough effective aperture area and good fittability to either certain satellite formats or rocket cargo bays. Indeed, a telescope of this type can be made nearly flat with relatively small Y side. Then it can be shortened in the Z direction using one or more tilting mirrors, in particular, a flat mirror located behind the primary, which redirects the beam along the primary’s longer side. In result, the larger sides of the telescope, $A_{x}$ and $A_{z}$, are essentially determined by the primary’s length $L_{x}$ and distance between the primary and secondary mirrors, correspondingly. In particular, such OS can be fitted within the 6U, 9U and other specific small satellite formats, where 1U is a standard CubeSat size measuring $10\times 10\times 10$ cm.
@@
To conclude, we presented a generic concept of an anamorphic telescope of a new type that contains cylindrical and toroidal optical surfaces, and can possess two large magnifications in the orthogonal directions of the image plane corresponding to the two independent focal distances, which allow to control both dimensions of the filed of view. The primary of this telescope can be contracted in one dimension keeping reasonably small aberrations and avoiding reduction of its field of view. This concept is promising for the sector of space optical industry, where the instrument’s shape and mass are of particular importance.
@@ Acknowledgments
The author thanks Barbara Pintaske and Daniela Ponce for their attention when providing the optical software of the SYNOPSYS, and the JATIS Reviewers for their useful comments.